Skip to end of metadata
Go to start of metadata

Unscheduled:

  • March 29th: The link to DDB from Kalix was down due to a fiber problem. The fiber provider has now fixed the problem and the service is restored.
  • March 25th: We were experiencing flaps on FPC0 on tug-r2.nordu.net router. The line card was replaced to fix the issue.
  • March 21st: The link between SMHI2 and LIU went down due to power outage in Norrköping.
  • March 16th: Link went down due to bad optical levels. Provider fixed the issue and the optical level are restored.
  • March 15th: The link between Örebro core router and ORU1 was down.
  • March 6th: The core router in Jönköping switched routing engine, RC unknown.
  • February 24th: The link between Svappavaara and Junosuando was down. Our provider fixed the issue and links are up. The OTDR and RFO are followed in SUNETTICKET-4202.

Scheduled:

  • March 31st: Two new switches were installed in Uppsala. The 1G ports on the router uu-r2 were migrated into the new switch uu-sw2.
    Migration of uu-br1 into the uu-sw1 will be done at a later stage.
  • March 29th: Fiber supplier performed fiber maintenance between Trollhättan and Göteborg.
  • March 29th: Faulty line card replaced on the router in Stockholm.
  • March 28th: SIPRI is moved over to a new link towards SUNET-C.
  • March 27th: Provider performed maintenance between the sites BOS-LJE2.
  • March 26th: Fiber supplier performed fiber maintenance between Abisko and Kiruna.
  • March 23rd: The maintenance was performed, and there was traffic between SUNET-C and SIPRI. However, the hosts at SIPRI were not able to reach any sites outside the SUNET network due to a most likely firewall issue. The traffic is diverted back to the old fiber and a new attempt will be made on a new date.
  • March 22nd: The fiber migration was cancelled and will be taken on a new date. SIPRI prefers to perform it after working hours.
  • March 17th: Supplier investigated the fiber between Sundsvall and MIUN1 for fault searching.
  • March 16th: Fiber provider performed maintenance in Stockholm affecting RKH and SH.
  • March 16th: Fiber provider performed maintenance between Karsefors (KFS) and Mörarp (MAP2)
  • March 15th: Supplier performed planned power works on an optical line between Kolbotten and Hedenlunda.
  • March 14th: The router in Trollhättan, hv-r1, was upgraded with a new software release to solve some issues.
  • March 14th: Fiber provider performed fiber maintenance between Borås and Ulricehamn
  • March 14th: The maintenance is over. We have lesser attenuation on both directions.
  • March 13th: The core router in Västerås was upgraded with a new software release.
  • March 9th: Supplier did fiber work on the stretch between Luleå and Piteå.
  • March 7th: A fiber work was carried out in Örebro.
  • March 6th: Fiber provider perform a planned maintenance on the stretch between Kiruna and Junosuando.
  • March 2nd: Maintenance adding the DVD site in the NMS is finished.
  • March 2nd: A new site was planned to be added between Norrköping and Hedenlunda but the work had to be cancelled due to wrong power cables shipped from equipment provider.

Scope

Unscheduled

Scheduled

Total

Hardware

1 pcs 00:02

2 pcs 00:10

3 pcs 00:12

Link

5 pcs 1d 11:10

13 pcs 1d 18:37

18 pcs 3d 05:47

None

1 pcs 00:11

5 pcs 13:46

6 pcs 13:57

Routing

0 pcs

0 pcs

0 pcs

Software

0 pcs

1 pcs 00:09

1 pcs 00:09

  • No labels